X=6, y=-4, a=3
(5x - 2y)(-2a)
evaluate.

Answer by stanbon(29482)

X=6, y=-4, a=3
(5x - 2y)(-2a)
---------------------
Substitute to get:
(5*6 - 2*-4)(-2*3)
-----
= (30+8)(-6)
---
= 38 *(-6)
---
= -228
=================
Cheers,
Stan H.